# Beam energy dependence of transverse momentum distribution and elliptic flow in Au-Au collisions using HYDJET++ model

### v1

@@ Abstract
In this work, we present the transverse momentum spectra and elliptic flow ($v_{2}$) of $\pi^{\pm},k^{\pm},$ $p$ and $\bar{p}$ in Au-Au collisions at $\sqrt{s_{NN}}=$ 62.4, 39.0, 27.0, 19.6 and 11.5 GeV using HYDJET++ model. For each beam energy the transverse momentum spectra are shown in four centrality classes at mid rapidity $|y|\leq$ 0.1 and for elliptic flow ($v_{2}$) we have shown results for minimum bias events. We have performed a qualitative comparison between the hard and soft processes at different beam energies. The HYDJET++ model calculations for particle ratios, $p_{T}$ spectra, and $v_{2}$ agree well with the available experimental data. The HYDJET++ parameters like baryonic and strangeness chemical potential, and chemical and thermal freeze-out temperature vary with collision energy. Azimuthal spatial anisotropy ($\epsilon$) and azimuthal momentum anisotropy ($\delta$) decrease with collision energy which leads to a smaller value of $v_{2}$ at lower collision energies. Furthermore, the hadrons containing strange quarks tend to have smaller values of spatial and momentum anisotropy than the non-strange hadrons.
@@ I Introduction
Heavy Ion collision experiments at RHIC and LHC have a common objective to study an exotic state of matter called Quark Gluon Plasma(QGP) formed at extreme temperature and density which is strong enough to de-confine the quarks and gluons [1, 2, 3]. One of the plausible approaches to study such a de-confined medium is to vary the collision energy and study its signals [4] like elliptic flow [5], jet quenching[6], strangeness abundance [7, 8], $j/\psi$ suppression [9], etc. Such an approach can be used for scanning the phase diagram of QCD as both temperature and chemical potential can be changed by varying collision energy [10, 11, 12]. Another advantage of using such a method is the possibility of finding the QCD critical point, which marks the end of the first-order phase transition [13, 14]. The critical point can be identified by enhanced fluctuations in conserved quantities like baryon number, charge, strangeness, etc [15, 16, 17].
@@
An experimental approach to study the collision systems at wide range of beam energies was made in the RHIC Beam Energy Scan(BES) program [18, 19, 20] which explored thirteen beam energies from the top RHIC energy of $\sqrt{s_{NN}}=$ 200 GeV to the lowest energy of 3 GeV. A series of experiments under the BES program tried to find the phase boundary and critical point by looking for the signals of QGP at different collision energies. The BES program generated a large volume of experimental data for different important observables like $p_{T}$ spectra, $v_{2}$, higher moments, etc for different collision energies. The highest energy of the BES program (i.e. 200 GeV) has been studied widely, while more phenomenological studies need to be performed for these smaller systems to study their behavior and mechanism of particle production. Furthermore, a qualitative difference between experimental data and theoretical model calculations or a deviation from the trend of model parameters at a particular beam energy could suggest a change in the particle production mechanism. Hence, the beam energy dependence of different signals and the model parameters need to be studied using a suitable theoretical model.
@@
Some theoretical attempts were made before to predict the beam energy dependence of signals like $v_{2}$ using models like AMPT, UrQMD, etc [21, 22]. A hydrodynamical treatment can be used to get a new insight into this problem. The theoretical model we chose to employ is the HYDJET++ model. It gives the option to study the collision systems as a combination of soft QCD processes (HYDrodynamics) and hard QCD processes (JETs). Our main goal is to study the bulk properties of the QGP medium produced in lower collision energies such as $p_{T}$ spectra, particle multiplicities, and elliptic flow, which are best described by the Monte Carlo event generators. HYDJET++ model has been successful in reproducing the experimental data at LHC energies and the top RHIC energy [23, 24, 25]. One of the main distinctions between the top RHIC energy, the LHC energies, and these lower energies is the large values of the chemical potential of these smaller systems. The model needs to be tested at lower collision energies to check the validity and implications of such treatment.
@@ II HYDJET++ MODEL
The hard part of the model is generated by PYQUEN. The initial patron spectra and jet production vertices from binary N-N sub-collisions are generated by PYTHIA 6 [27]. The jets produced in Au-Au collisions are a combination of jets produced from these binary sub-collisions. PYQUEN calculates the rescattering path of partons in the dense QGP medium and the radiative and collisional energy loss. The final hadronization is done according to the Lund string model. The nuclear shadowing for each impact parameter is done according to Glauber-Gribov theory [28].
@@
where, $\mu_{\pi^{+}}^{eff}$ is the effective chemical potential of $\pi^{+}$ at thermal freeze-out. The values of $T_{th}$ and $\mu_{\pi^{+}}^{eff}$ are chosen to get the best match with $p_{T}$ distribution measured in RHIC BES [19]. For this work, we have assumed $T_{th}$ and $\mu_{\pi^{+}}^{eff}$ to be independent of centrality. Assuming Boltzmann approximation in eq.  (2) for particles heavier than pions [31], from eq.  (2) and eq.  (4) the chemical potential of $i^{th}$ hadron species at thermal freeze-out:
@@
$u_{\perp}=\gamma v_{\perp}=cosh\eta^{u}\gamma_{\perp}v_{\perp}$, $\gamma_{\perp}=cosh\rho^{u}$. Due to the presence of non-zero azimuthal anisotropy, the azimuthal angle of the fluid velocity vector is not identical to the spatial azimuthal angle i.e.
@@
The transverse flow rapidity $\rho^{u}$ is related to $\rho^{{}^{\prime}u}$ as follows:
@@
Elliptic flow is a result of pressure gradient in the overlap region. The expansion would be stronger along the reaction plane where the pressure gradient is larger. Due to this expansion, the initial spatial anisotropy keeps changing [32]. Since the model does not trace the evolution of spatial anisotropy, we have only used the spatial anisotropy at thermal freeze-out. The initial spatial anisotropy depends on the impact parameter and nuclear geometry. Hence we have assumed a simple scaling between the eccentricity of the elliptical overlap region($\epsilon_{0}=b/2R_{A}$) and the spatial anisotropy at thermal freeze-out i.e.
@@
The parameter ”k” is a scaling factor that indicates the change in azimuthal spatial anisotropy during evolution. The parameters $\delta$ and $\epsilon$ are related to each other by their dependence $v_{2}$, an empirical relation between the parameters estimated by the hydro-dynamical approach in the corresponding paper [5]:
@@
Using the fact that $v_{2}\propto\epsilon_{0}$ and $\epsilon\propto\epsilon_{0}$ the relation between $\epsilon$ and $\delta$ can be calculated using eq.  (12),
@@
In this work, we have fixed the value of $\mathrm{C=2}$, and the value of scaling factor $\mathrm{k}$ is chosen to get the best match with the experimental data.
@@
Besides the above-synchronized formalism, the HYDJET++ model also has certain limitations. HYDJET++ model gives good results for the central and semi-central collision events but it fails to give convincing results for the peripheral collisions. The model can only be used for symmetric collisions (i.e. both the colliding nuclei should be of the same element). It works reliably for relatively heavier nuclei with nuclear mass A$>$40. The model best describes the experimental data in the mid-rapidity region. Another limitation of the HYDJET++ model is that it should be used for $\sqrt{s_{NN}}>10$ GeV. For the limitations described above we have chosen to limit our results up to 11.5 $\mathrm{GeV}$ and presented all our results in the mid-rapidity region.
@@ III RESULTS AND DISCUSSIONS
We have simulated Au-Au collisions at a center of mass energies 62.4 GeV, 39.0 GeV, 27.0 GeV, 19.6 GeV, and 11.5 GeV using the HYDJET++ event generator and compared our results with the experimental data from STAR and PHENIX detectors. The kinematic range of our results is kept the same as the experimental data. The rapidity range for particle ratios and $p_{T}$ spectra is $|y|\leq 0.1$ and for $v_{2}$ rapidity range is $|y|\leq$ 0.35 for PHENIX data and $|y|\leq$ 1.0 for STAR data. The exact number of events can be found in Table 1.
@@
Figure 6: Transverse momentum spectra for different centralities at $\sqrt{s_{NN}}$= 62.4 GeV. Markers represent experimental data and HYDJET++ results are shown as lines.
Figure 7: Transverse momentum spectra for different centralities at $\sqrt{s_{NN}}$= 39.0 GeV. Markers represent experimental data and HYDJET++ results are shown as lines.
Figure 8: Transverse momentum spectra for different centralities at $\sqrt{s_{NN}}$= 27.0 GeV. Markers represent experimental data and HYDJET++ results are shown as lines.
Figure 9: Transverse momentum spectra for different centralities at $\sqrt{s_{NN}}$= 19.6 GeV. Markers represent experimental data and HYDJET++ results are shown as lines.
Figure 10: Transverse momentum spectra for different centralities at $\sqrt{s_{NN}}$= 11.5 GeV. Markers represent experimental data and HYDJET++ results are shown as lines.
@@ III.1 Particle Ratios
Fig. 1 shows HYDJET++ results for $\pi^{-}/\pi^{+}$,$k^{-}/k^{+}$,$k^{+}/\pi^{+}$,$k^{-}/\pi^{-}$ and $\bar{p}/p$ ratios at various center of mass energies for most central collisions at mid rapidity. The $\pi^{-}/\pi^{+}$ ratio increases at lower collision energies but still remains close to 1. The $k^{-}/k^{+}$ ratio increases with beam energy and approaches unity at $\sqrt{s_{NN}}=62.4$ GeV. The $\bar{p}/p$ ratio also increases with beam energy, but the increase is more rapid than $k^{-}/k^{+}$. The $k^{+}/\pi^{+}$ ratios decrease with increasing beam energy while the $k^{-}/\pi^{-}$ ratio shows the opposite trend. Both these ratios become the same after 62.4 GeV due to smaller values of $\mu_{s}$ at higher collision energies. The model calculations agree well with the experimental data with a slight deviation at 11.5 GeV.
@@ III.2 $p_{T}$ spectra
In the HYDJET++ framework, the particles are produced as a combination of a hydro part and jets. So, It is crucial to check the contribution from both processes. Fig. 2 shows the contributions from hydro and jets in $p_{T}$ spectra of $\pi^{\pm}$ for most central collisions for $\sqrt{s_{NN}}$=62.4, 27.0 and 11.5 GeV. It is evident that the hydro part predominates the $p_{T}$ spectra at low $p_{T}$, and the contribution from jets starts to affect the overall form of $p_{T}$ spectra at relatively higher $p_{T}$. It’s worth noting that the contributions from both processes drop with decreasing beam energy. However, the contribution from jets decreases more rapidly than that of the hydro part. In Fig.  3 we have shown the effect of both processes for $k^{\pm}$ , the contributions are similar to that of charged pions, however for $k^{-}$ the contributions from the soft part decrease more rapidly with decreasing beam energy than $k^{+}$. The effects from hydro and jets for $p$ and $\bar{p}$ are shown in Fig. 4. In the $p_{T}$ spectra of protons, the contributions from jets decrease with decreasing beam energy while the contributions from the hydro part increase rapidly. As a result the overall contribution increases. This can be attributed to an increase in baryon density at lower collision energies [19]. For $\bar{p}$ contributions from both processes decrease rapidly with decreasing beam energy. We have discussed the difference in $p_{T}$ of particles and antiparticles later. A similar analysis for Au-Au collisions at 200 GeV using the HYDJET++ model for kaons and pions can be found in the corresponding papers [25, 24].
@@
In Fig. 6 to Fig. 10 we have shown the transverse momentum spectra at low $p_{T}$ for different beam energies. Our results are shown in four centrality classes i.e. (0-5)%, (10-20)%, (30-40)%, and (50-60)%. For comparison, we have shown the experimental data from STAR[19, 36]. The $p_{T}$ spectra of $\pi^{\pm}$ and $k^{\pm}$ show a good match with the experimental data in central and semi-central collision while the model slightly over-predict $k^{\pm}$ at peripheral collisions for lower beam energies. The model under-predicts the $p_{T}$ spectra of $p$ and $\bar{p}$ up to $p_{T}=0.5$ GeV$/c$ after 0.5 GeV$/c$ there is a good match. The results for peripheral collisions at lower beam energy show a clear mismatch with the experimental data. This is due to a strong centrality dependence of $\mu_{B}$ which is observed at 11.5 and 19.6 GeV [19].
The invariant yield tends to decrease from central to peripheral collisions. This behavior is the same for all particles. There is no beam energy dependence of variation of $p_{T}$ spectra with collision centrality observed in our model calculations. The centrality dependence at lower collision energies is similar to the one observed for the top RHIC energy as well as LHC energies [37, 38].

### v2

@@ Abstract
In this work, we present the particle ratios, transverse momentum spectra, and elliptic flow ($v_{2}$) of $\pi^{\pm},k^{\pm},$ $p$, and $\bar{p}$ in Au-Au collisions at $\sqrt{s_{NN}}=$ 62.4, 39.0, 27.0, 19.6 and 11.5 GeV using HYDJET++ model. The particle ratios match the experimental data that validates the Cleymans-Reidlich parameterization of freeze-out parameters at lower beam energies under the HYDJET++ framework. The lower collision energies produce a system of high baryon chemical potential ($\mu_{B}$) and have a lower inelastic cross section. The interplay between these effects affects the overall shape of the $p_{T}$ spectra. The HYDJET++ model calculations for $p_{T}$ spectra agree well with the available experimental data. The invariant yield ratio of central and peripheral collisions is independent of beam energy. The elliptic flow is calculated based on the scaling between initial and final azimuthal spatial anisotropy ($k$). This interpretation of $v_{2}$ successfully describes the experimental data for all the collision energies studied in this work. The positive correlation of $k$ with beam energy leads to a small $v_{2}$ at lower collision energies. The hadrons containing strange quarks tend to have smaller values of $k$ than the non-strange hadrons.
@@ I Introduction
Heavy Ion collision experiments at RHIC and LHC have a common objective to study an exotic state of matter called Quark Gluon Plasma(QGP) formed at extreme temperature and density, which is strong enough to de-confine the quarks and gluons [1, 2, 3]. One of the plausible approaches to study such a de-confined medium is to vary the collision energy and study its signals [4] like elliptic flow [5], jet quenching[6], strangeness abundance [7, 8], $j/\psi$ suppression [9], etc. Such an approach can be used for scanning the phase diagram of QCD as both temperature and chemical potential can be changed by varying collision energy [10, 11, 12]. Another advantage of using such a method is the possibility of finding the QCD critical point, which marks the end of the first-order phase transition [13, 14]. The critical point can be identified by enhanced fluctuations in conserved quantities like baryon number, charge, strangeness, etc [15, 16, 17].
@@
An experimental approach to study the collision systems at wide range of beam energies was made in the RHIC Beam Energy Scan(BES) program [18, 19, 20] which explored thirteen beam energies from the top RHIC energy of $\sqrt{s_{NN}}=$ 200 GeV to the lowest energy of 3 GeV. A series of experiments under the BES program tried to find the phase boundary and critical point by looking for the signals of QGP at different collision energies. The BES program generated a large volume of experimental data for different important observables like $p_{T}$ spectra, $v_{2}$, higher moments, etc, for different collision energies. The highest energy of the BES program (i.e. 200 GeV) has been studied widely, while more phenomenological studies need to be performed for these smaller systems to study their behavior and mechanism of particle production. Furthermore, a qualitative difference between experimental data and theoretical model calculations or a deviation from the trend of model parameters at a particular beam energy could suggest a change in the particle production mechanism. Hence, the beam energy dependence of different signals and the model parameters need to be studied using a suitable theoretical model.
@@
Some theoretical attempts were made before to predict the beam energy dependence of signals like $v_{2}$ using models like AMPT, UrQMD, etc [21, 22]. A hydrodynamical treatment can be used to get a new insight into this problem. The theoretical model we chose to employ is the HYDJET++ model. It gives the option to study the collision systems as a combination of soft QCD processes (HYDrodynamics) and hard QCD processes (JETs). Our main goal is to study the bulk properties of the QGP medium produced in lower collision energies such as $p_{T}$ spectra, particle multiplicities, and elliptic flow, which are best described by the Monte Carlo event generators. The HYDJET++ model has been successful in reproducing the experimental data at LHC energies and the top RHIC energy [23, 24, 25]. One of the main distinctions between the top RHIC energy, the LHC energies, and these lower energies is the large values of the chemical potential of these smaller systems. The model needs to be tested at lower collision energies to check the validity and implications of such treatment.
@@ II HYDJET++ MODEL
The hard part of the model is generated by PYQUEN. The initial patron spectra and jet production vertices from binary N-N sub-collisions are generated by PYTHIA 6 [27]. The jets produced in Au-Au collisions are a combination of jets produced from these binary sub-collisions. PYQUEN calculates the rescattering path of partons in the dense QGP medium and the radiative and collisional energy loss. The final hadronization is done according to the Lund string model. The nuclear shadowing for each impact parameter is done according to the Glauber-Gribov theory [28].
@@
where, $\mu_{\pi^{+}}^{eff}$ is the effective chemical potential of $\pi^{+}$ at thermal freeze-out. The values of $T_{th}$ and $\mu_{\pi^{+}}^{eff}$ are chosen to get the best match with $p_{T}$ distribution measured in RHIC BES [19]. Assuming Boltzmann approximation in eq.  (2) for particles heavier than pions [31], from eq.  (2) and eq.  (4) the chemical potential of $i^{th}$ hadron species at thermal freeze-out:
@@
$u_{\perp}=\gamma v_{\perp}=cosh\eta^{u}\gamma_{\perp}v_{\perp}$, $\gamma_{\perp}=cosh\rho^{u}$. Due to the presence of non-zero azimuthal anisotropy, the azimuthal angle of the fluid velocity vector is not identical to the spatial azimuthal angle, i.e.
@@
The parameter $\delta(b)$ is the azimuthal momentum anisotropy, a more detailed description of $\delta(b)$ is given later in the text. The transverse flow rapidity $\rho^{u}$ is related to $\rho^{{}^{\prime}u}$ as follows:
@@
Elliptic flow is a result of pressure gradient in the overlap region. The expansion would be stronger along the reaction plane where the pressure gradient is larger. Due to this expansion, the initial spatial anisotropy keeps changing [32]. Since the model does not trace the evolution of spatial anisotropy, we have only used the spatial anisotropy at thermal freeze-out. The initial spatial anisotropy depends on the impact parameter and nuclear geometry. Hence, we have assumed a simple scaling between the eccentricity of the elliptical overlap region($\epsilon_{0}=b/2R_{A}$) and the spatial anisotropy at thermal freeze-out, i.e.
@@
The parameter “$k$” is a scaling factor that indicates the change in azimuthal spatial anisotropy during evolution. It is a common practice to treat the parameters $\epsilon$ and $\delta$ as free parameters, but in such a treatment, the parameters lose all relation with the initial collision geometry. The relation can be established through the scaling factor “$k$”. The scaling factor relates the eccentricity of the overlap region $\epsilon_{0}$ with the azimuthal spatial anisotropy of the fireball at freeze-out $\epsilon$, which then determines the momentum anisotropy $\delta$. Hence, the scaling factor “$k$” determines both the spatial and momentum anisotropy parameters and the elliptic flow of final state hadrons. An empirical relation between the parameters estimated by the hydro-dynamical approach in the corresponding paper [5]:
@@
Using the fact that $v_{2}\propto\epsilon_{0}$ and $\epsilon\propto\epsilon_{0}$, the relation between $\epsilon$ and $\delta$ can be calculated using eq.  (12),
@@
The conversion of spatial anisotropy to momentum anisotropy, which is controlled by the parameter “C”, is assumed to be constant across all collision energies. In this work, we have fixed the value of $\mathrm{C=2}$, which was used at 200 GeV in the corresponding paper [24]. The value of the scaling factor “$k$” is chosen to get the best match with the experimental data.
@@
Besides the above-synchronized formalism, the HYDJET++ model also has certain limitations. HYDJET++ model gives good results for the central and semi-central collision events, but it fails to give convincing results for the peripheral collisions. The model can only be used for symmetric collisions (i.e., both the colliding nuclei should be of the same element). It works reliably for relatively heavier nuclei with nuclear mass A$>$40. The model best describes the experimental data in the mid-rapidity region. Another limitation of the HYDJET++ model is that it should be used for $\sqrt{s_{NN}}>10$ GeV. For the limitations described above, we have chosen to limit our results up to 11.5 $\mathrm{GeV}$ and presented all our results in the mid-rapidity region.
@@ III RESULTS AND DISCUSSIONS
We have simulated Au-Au collisions at center-of-mass energies 62.4 GeV, 39.0 GeV, 27.0 GeV, 19.6 GeV, and 11.5 GeV using the HYDJET++ event generator and compared our results with the experimental data from STAR and PHENIX detectors. The kinematic range of our results is kept the same as the experimental data. The rapidity range for particle ratios and $p_{T}$ spectra is $|y|\leq 0.1$, and for $v_{2}$ rapidity range is $|y|\leq$ 0.35 for PHENIX data and $|y|\leq$ 1.0 for STAR data. The exact number of events can be found in Table 1.
@@
Figure 6: Transverse momentum spectra for different centralities at $\sqrt{s_{NN}}$= 62.4 GeV. Markers represent experimental data, and HYDJET++ results are shown as lines.
Figure 7: Transverse momentum spectra for different centralities at $\sqrt{s_{NN}}$= 39.0 GeV. Markers represent experimental data, and HYDJET++ results are shown as lines.
Figure 8: Transverse momentum spectra for different centralities at $\sqrt{s_{NN}}$= 27.0 GeV. Markers represent experimental data, and HYDJET++ results are shown as lines.
Figure 9: Transverse momentum spectra for different centralities at $\sqrt{s_{NN}}$= 19.6 GeV. Markers represent experimental data, and HYDJET++ results are shown as lines.
Figure 10: Transverse momentum spectra for different centralities at $\sqrt{s_{NN}}$= 11.5 GeV. Markers represent experimental data, and HYDJET++ results are shown as lines.
@@ III.1 Particle Ratios
Fig. 1 shows HYDJET++ results for $\pi^{-}/\pi^{+}$,$k^{-}/k^{+}$,$k^{+}/\pi^{+}$,$k^{-}/\pi^{-}$ and $\bar{p}/p$ ratios at various center of mass energies for most central collisions at mid rapidity. The $\pi^{-}/\pi^{+}$ ratio increases at lower collision energies but still remains close to 1. The $k^{-}/k^{+}$ ratio increases with beam energy and approaches unity at $\sqrt{s_{NN}}=62.4$ GeV. The $\bar{p}/p$ ratio also increases with beam energy, but the increase is more rapid than $k^{-}/k^{+}$. The $k^{+}/\pi^{+}$ ratios decrease with increasing beam energy, while the $k^{-}/\pi^{-}$ ratio shows the opposite trend. Both these ratios become the same after 62.4 GeV due to smaller values of $\mu_{s}$ at higher collision energies. The model calculations agree well with the experimental data with a slight deviation at 11.5 GeV.
Figure 17: Relation between scaling factor k and beam energy. The lines show the fitting function from eq.  (15).
Figure 18: $v_{2}$ of identified hadrons at 62.4 GeV with different $k$ values. The red lines represent the results with the values of $k$ used for non-strange hadrons, and the blue lines represent the results with the values used for strange hadrons.
@@ III.2 $p_{T}$ spectra
In the HYDJET++ framework, the particles are produced as a combination of a hydro part and jets. So, it is crucial to check the contribution from both processes. Fig. 2 shows the contributions from hydro and jets in $p_{T}$ spectra of $\pi^{\pm}$ for most central collisions for $\sqrt{s_{NN}}$=62.4, 27.0 and 11.5 GeV. It is evident that the hydro part predominates the $p_{T}$ spectra at low $p_{T}$, and the contribution from jets starts to affect the overall form of $p_{T}$ spectra at relatively higher $p_{T}$. It’s worth noting that the contributions from both processes drop with decreasing beam energy. However, the contribution from jets decreases more rapidly than that of the hydro part. In Fig.  3 we have shown the effect of both processes for $k^{\pm}$ , the contributions are similar to that of charged pions, however for $k^{-}$, the contributions from the soft part decrease more rapidly with decreasing beam energy than $k^{+}$. The effects from hydro and jets for $p$ and $\bar{p}$ are shown in Fig. 4. In the $p_{T}$ spectra of protons, the contributions from jets decrease with decreasing beam energy while the contributions from the hydro part increase rapidly. As a result, the overall contribution increases. This can be attributed to an increase in baryon density at lower collision energies [19]. For $\bar{p}$, contributions from both processes decrease rapidly with decreasing beam energy. However, the hydro contribution decreases faster than the jets, leading to the modification of low $p_{T}$ spectra by jets. We have discussed the difference in $p_{T}$ of particles and antiparticles later. A similar analysis for Au-Au collisions at 200 GeV using the HYDJET++ model for kaons and pions can be found in the corresponding papers [25, 24].
@@
In Fig. 6 to Fig. 10, we have shown the transverse momentum spectra at low $p_{T}$ for different beam energies. Our results are shown in four centrality classes i.e. (0-5)%, (10-20)%, (30-40)%, and (50-60)%. For comparison, we have shown the experimental data from STAR[19, 36]. The $p_{T}$ spectra of $\pi^{\pm}$ and $k^{\pm}$ show a good match with the experimental data in central and semi-central collision while the model slightly over-predict $k^{\pm}$ at peripheral collisions for lower beam energies. The model under-predicts the $p_{T}$ spectra of $p$ and $\bar{p}$ up to $p_{T}=0.5$ GeV$/c$, and after 0.5 GeV$/c$ there is a good match. The results for peripheral collisions at lower beam energy show a clear mismatch with the experimental data. This is due to a strong centrality dependence of $\mu_{B}$, which is observed at 11.5 and 19.6 GeV [19].
The centrality dependence at lower collision energies is similar to the one observed for the top RHIC energy as well as LHC energies [37, 38]. The invariant yield tends to decrease from central to peripheral collisions. This behavior is the same for all primary charged hadrons. The ratio of the invariant yields of central and peripheral collision is the same for all collision energies for all particles. Thus, there is no beam energy dependence of variation of $p_{T}$ spectra with collision centrality observed in our model calculations.
Recent experiments have shown that high $p_{T}$ suppression due to jet quenching is not observed for lower collision energies, instead an enhancement is observed, possibly due to the Cronin effect [39]. Since the HYDJET++ model does not have any mechanism for enhancement of high $p_{T}$ hadrons the results will underpredict data. Hence, we have only shown our results at low $p_{T}$ where no such effects are observed.